# One Prompt Fits All: Universal Graph Adaptation for Pretrained Models

### v1

@@ B.9 Hyperparameter Settings
PubMed 0.0005 0.05 1 0.9999 0.0001 0.05 50 0.9999 0.0005 0.05 10 0.99999
Cornell 0.0001 0.005 50 0.9999 0.001 0.01 50 0.9999 0.00005 0.001 50 0.9999
Texas 0.01 0.005 50 0.9999 0.0001 0.00005 50 0.9999 0.00001 0.0001 50 0.9999
Wisconsin 0.0001 0.0005 50 0.9999 0.005 0.0001 50 0.9999 0.00001 0.0001 50 0.9999
Chameleon 0.01 0.001 50 0.9999 0.00001 0.05 10 0.999 0.00001 0.05 10 0.999
Actor 0.00005 0.005 50 0.9999 0.00001 0.01 50 0.9999 0.0005 0.005 50 0.9999
Squirrel 0.0005 0.05 50 0.9999 0.0005 0.01 50 0.99 0.0001 0.0001 50 0.9999
@@
Wisconsin 0.001 0.00005 50 0.9999 0.0001 0.0005 50 0.999 0.0001 0.0001 50 0.9999
Chameleon 0.01 0.0005 1 0.99999 0.001 0.001 50 0.9999 0.005 0.05 50 0.99999
Actor 0.00001 0.01 50 0.9999 0.005 0.01 50 0.9999 0.005 0.05 50 0.9999
@@ Appendix C Related Works
Graph Pretraining. Graph pretraining has emerged as a powerful paradigm for learning generalizable and transferable representations from large-scale unlabeled graph data, aiming to mitigate the dependency on labeled data in downstream tasks. Unlike traditional supervised methods that require extensive manual annotations [3, 4, 48], graph pretraining leverages self-supervised strategies to capture structural and semantic patterns in graphs. Graph Self-Supervised Learning (GSSL) [49] currently has attracted widespread attention in the graph community, which mainly designs self-supervised objective functions to train the model based on maximizing Mutual Information (MI). As a classic paradigm, DGI [28] maximizes mutual information between node representations and the summary of the graph. GGD [50] further explores the DGI, summarizing it as a group discrimination task, greatly reducing the computational time overhead. MVGRL [51] introduces graph diffusion to generate different scale subgraphs to improve the pipeline of DGI. GRACE [29] uses InfoNCE [45, 46] to optimize by maximizing the similarity of two augmented nodes generated by the same node and minimizing the similarity of other nodes to train the model. GCA [47] improve the augmentation strategy by defining the importance of different nodes and edges to preserve semantic information. GraphCL [37] and JOAO [52] use this paradigm to global-global graph representation contrast. AD-GCL [53] has similar idea, which designs a learnable augmentation strategy, and trains the model by maximizing mutual information between node representation and augmented graph. Some works [54, 55, 56, 57, 58] focus on design effective sampling strategy, and other works [59, 60] further introduce additional knowledge into GCL. AFGRL [61], SimGRACE [62], NeCo [63] and AFGCL [64] propose augmentation free paradigm to optimize sampling. PolyGCL [65] and S3GCL [66] focus on designing polynomials with learnable filters to generate different spectral contrastive views. GraphMAE [30], as the method of masked graph autoencoders, utilizes randomly mask mechanism from the input with the graph autoencoder [67] architecture to reconstruct the node features or structures. GraphMAE2 [68] proposes multiview random re-mask decoding, the node representations are randomly re-masked multiple times, to introduce randomness into feature reconstruction. BGRL [31] adopts BYOL [69], which trains the online encoder by predicting the target encoder to generate efficient node representations. This backbone is followed by some recent works, such as SGRL [70] and SGCL [32].
@@ Appendix E Broader Impacts
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [Yes]
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [Yes]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: [Yes]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  54. Justification: This item is not relevant to our work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [N/A]
  64. Justification: This item is not relevant to our work.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  69. Justification: This item is not relevant to our work.
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  74. Justification: This item is not relevant to our work.
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  80. Guidelines:

### v2

@@ 8 Acknowledgments
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [Yes]
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [Yes]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: [Yes]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  54. Justification: This item is not relevant to our work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [N/A]
  64. Justification: This item is not relevant to our work.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  69. Justification: This item is not relevant to our work.
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  74. Justification: This item is not relevant to our work.
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  80. Guidelines:
@@ B.9 Hyperparameter Settings
PubMed 0.0005 0.001 1 0.9999 0.0001 0.05 50 0.9999 0.0005 0.05 10 0.99999
Cornell 0.001 0.0005 50 0.99 0.001 0.01 50 0.9999 0.00005 0.001 50 0.9999
Texas 0.00001 0.0001 50 0.999 0.0001 0.00005 50 0.9999 0.00001 0.0001 50 0.9999
Wisconsin 0.0001 0.001 50 0.999 0.005 0.0001 50 0.9999 0.00001 0.0001 50 0.9999
Chameleon 0.00005 0.001 10 0.9999 0.00001 0.05 10 0.999 0.00001 0.05 10 0.999
Actor 0.001 0.01 50 0.999 0.00001 0.01 50 0.9999 0.0005 0.005 50 0.9999
Squirrel 0.00005 0.005 50 0.99999 0.0005 0.01 50 0.99 0.0001 0.0001 50 0.9999
@@
Wisconsin 0.0001 0.01 50 0.999 0.0001 0.0005 50 0.999 0.0001 0.0001 50 0.9999
Chameleon 0.005 0.001 1 0.99999 0.001 0.001 50 0.9999 0.005 0.05 50 0.99999
Actor 0.0005 0.01 50 0.9999 0.005 0.01 50 0.9999 0.005 0.05 50 0.9999
@@ Appendix C Related Works
Graph Pretraining. Graph pretraining has emerged as a powerful paradigm for learning generalizable and transferable representations from large-scale unlabeled graph data, aiming to mitigate the dependency on labeled data in downstream tasks. Unlike traditional supervised methods that require extensive manual annotations [3, 4, 48], graph pretraining leverages self-supervised strategies to capture structural and semantic patterns in graphs. Graph Self-Supervised Learning (GSSL) [49] currently has attracted widespread attention in the graph community, which mainly designs self-supervised objective functions to train the model based on maximizing Mutual Information (MI). As a classic paradigm, DGI [28] maximizes mutual information between node representations and the summary of the graph. GGD [50] further explores the DGI, summarizing it as a group discrimination task, greatly reducing the computational time overhead. MVGRL [51] introduces graph diffusion to generate different scale subgraphs to improve the pipeline of DGI. GRACE [29] uses InfoNCE [45, 46] to optimize by maximizing the similarity of two augmented nodes generated by the same node and minimizing the similarity of other nodes to train the model. GCA [47] improve the augmentation strategy by defining the importance of different nodes and edges to preserve semantic information. GraphCL [37] and JOAO [52] use this paradigm to global-global graph representation contrast. AD-GCL [53] has similar idea, which designs a learnable augmentation strategy, and trains the model by maximizing mutual information between node representation and augmented graph. Some works [54, 55, 56, 57, 58, 59, 60, 61] focus on design effective sampling strategy, and other works [62, 63] further introduce additional knowledge into GCL. AFGRL [64], SimGRACE [65], NeCo [66] and AFGCL [67] propose augmentation free paradigm to optimize sampling. PolyGCL [68] and S3GCL [69] focus on designing polynomials with learnable filters to generate different spectral contrastive views. GraphMAE [30], as the method of masked graph autoencoders, utilizes randomly mask mechanism from the input with the graph autoencoder [70] architecture to reconstruct the node features or structures. GraphMAE2 [71] proposes multiview random re-mask decoding, the node representations are randomly re-masked multiple times, to introduce randomness into feature reconstruction. BGRL [31] adopts BYOL [72], which trains the online encoder by predicting the target encoder to generate efficient node representations. This backbone is followed by some recent works, such as SGRL [73] and SGCL [32].